# A Survey of Backdoor Attacks and Defenses on Large Language Models: Implications for Security Measures

### v2

@@ 3.1 Backdoor Attack based on Full-parameter Fine-tuning
The efficiency of LLMs has been proven in various NLP tasks, demonstrating their ability to understand and generate text in ways that are both sophisticated and contextually relevant. These models have become indispensable tools in machine translation (Zhang et al.,, 2023; Garcia et al.,, 2023), summary generation (Nguyen et al.,, 2021; Nguyen and Luu,, 2022; Zhao et al.,, 2022; Zhao et al., 2023a, ), and recommendation systems (Ma et al.,, 2016; Li et al., 2024a, ). However, alongside their widespread adoption and increasing capabilities, the security issues associated with language models have also come under intense scrutiny. Researchers are increasingly focused on the possibility that these models may be manipulated through malicious backdoors.
@@
You et al., (2023) introduce a backdoor attack algorithm, named LLMBkd, which leverages LLMs to automatically embed a specified textual style as a trigger within samples. Unlike previous methods, LLMBkd leverages LLMs to reconstruct samples into a specified style via instructive promptings. Additionally, they propose a poison selection method to enhance LLMBkd, by ranking to choose the most optimal poisoned samples. Kandpal et al., (2023) explore the security of LLMs based on in-context learning. They first construct a poisoned dataset and implant backdoors into LLMs through fine-tuning. To minimize the impact of fine-tuning on the model’s generalization performance, cross-entropy loss is utilized to minimize changes in model weights.
@@
Shi et al., (2023) construct BadGPT, the first backdoor attack against reinforcement learning fine-tuning in LLMs. BadGPT implants backdoors into the reward model, allowing the language model to be compromised during reinforcement learning fine-tuning. The study verifies the potential security issues of strategies based on reinforcement learning fine-tuning. Wang et al., 2023b explore the potential security issues of RLHF, where attackers manipulate ranking scores by altering the rankings of any malicious text, leading to adversarially guided responses from LLMs. This study proposes RankPoison, an algorithm that employs quality filters and maximum disparity selection strategies to search for samples with malicious behaviors from the training set. Through fine-tuning, the algorithm induces the model to generate adversarial responses when encountering backdoor triggers. Li et al., 2023a utilize black-box generative models, such as ChatGPT, as a backdoor attack tool to construct the BGMAttack algorithm. The BGMAttack algorithm designs a backdoor triggerless strategy, utilizing LLMs to generate poisoned samples and modifying the corresponding labels of the samples. Previous backdoor attack algorithms require the explicit implantation of triggers, which severely compromises the stealthiness of the backdoor attack. Zhao et al., 2023b employ manually written prompt as trigger, obviating the need for implanting additional triggers and preserving the integrity of the training samples, enhancing the stealthiness of the backdoor attack. Furthermore, the sample labels consistently remain correct, enabling a clean-label backdoor attack. Tan et al., (2023) propose a more flexible backdoor attack algorithm, named TARGET, which utilizes GPT-4 as a backdoor attack tool to generate malicious templates that act as triggers. Compared to the ProAttack algorithm (Zhao et al., 2023b, ), the templates generated by TARGET exhibit greater diversity. Qi et al., (2023) validate the fragility of the safety alignment of LLMs across three dimensions. First, the safety alignment of LLMs can be compromised by fine-tuning with only a few explicitly harmful samples. Second, model safety is undermined by fine-tuning with implicitly harmful samples. Finally, under the influence of "catastrophic forgetting" (Kirkpatrick et al.,, 2017; Luo et al.,, 2023), model safety still significantly deteriorates even when fine-tuning on the original dataset. Li et al., 2024b introduce the BadEdit backdoor attack framework, which directly modifies a small number of LLM parameters to efficiently implement backdoor attacks while preserving model performance. Specifically, the backdoor injection problem is redefined as a knowledge editing problem. Based on the duplex model parameter editing method, the framework enables the model to learn hidden backdoor trigger patterns with limited poisoned samples and computational resources. Zou et al., (2024) explore the security of retrieval-augmented generation (RAG) in LLMs. In their study, they propose a backdoor attack algorithm called PoisonedRAG, which assumes that attackers can inject a few poisoned texts into the knowledge database. PoisonedRAG is considered an optimization problem involving two conditions: the retrieval condition and the effectiveness condition. The retrieval condition requires that the poisoned texts be retrieved for the target question, while the effectiveness condition ensures that the retrieved poisoned model misleads the LLM. Yang et al., (2024) investigate the security of LLM-based agents when faced with backdoor attacks. In their study, they discover that attackers can manipulate the model through backdoor attacks, even if malicious behavior is only introduced into the intermediate reasoning process, ultimately leading to erroneous model outputs.
@@
Notes Existing work has demonstrated that language models are susceptible to manipulation through backdoors. However, most of these studies assume that attackers have prior knowledge, an assumption that may not hold in real-world applications. Therefore, exploring task-agnostic backdoor attack algorithms is an issue that deserves continuous scrutiny. Furthermore, the full-parameter fine-tuning strategy also introduces additional overhead to the deployment of backdoor attacks.
@@ 3.2 Backdoor Attack based on Parameter-Efficient Fine-Tuning
Figure 2: Overview of the backdoor attack based on PEFT, where the fine-tuning algorithm employs prompt-tuning.
@@
Prompt-tuning Xue et al., (2024) introduce TrojLLM, a black-box framework that includes the trigger discovery algorithm and the progressive Trojan poisoning algorithm, capable of autonomously generating triggers with universality and stealthiness. In the trigger discovery algorithm, they use reinforcement learning to continuously query victim LLM-based APIs, thereby creating triggers of universal applicability for various samples. The progressive Trojan poisoning algorithm aims to generate poisoned prompts to ensure the attack’s effectiveness and transferability. Yao et al., (2024) introduce a novel two-stage optimization backdoor attack algorithm that successfully compromises both hard and soft prompt-based LLMs. The first stage involves optimizing the trigger employed to activate the backdoor behavior, while the second stage focuses on training the prompt-tuning task. Huang et al., 2023a propose a composite backdoor attack algorithm with enhanced stealth, named CBA. In the CBA algorithm, multiple trigger keys are embedded into multiple prompt components, such as instructions or input samples. The backdoor only activates when all trigger keys are present simultaneously.
@@
Notes The effectiveness of backdoor attacks, particularly those that target PEFT methods, has been clearly demonstrated. However, existing work primarily focuses on classification tasks. It is worth mentioning that extending these backdoor attacks to generative tasks, while simultaneously exploring clean-label backdoor attacks, presents a more significant challenge.
@@ 3.3 Backdoor Attack without Fine-tuning
Figure 3: Overview of the backdoor attack without fine-tuning, illustrating attacks on instructions and in-context learning.
LoRA In share-and-play settings, Liu et al., (2024) assume that the LoRA (Hu et al.,, 2021) algorithm could be a potential attacker capable of injecting backdoors into LLMs. They combine an adversarial LoRA with a benign LoRA to investigate attack methods that do not require backdoor fine-tuning. Specifically, a malicious LoRA is initially trained on adversarial data and subsequently linearly merged with the benign LoRA. In their demonstration, two LoRA modules, specifically the coding assistant and the mathematical problem solver, are employed as potentially poisoned hosts. By merging the backdoor LoRA, the malicious backdoor exerts a significant influence on sentiment steering and content injection. Although the experiments demonstrate that LoRA modules can serve as potential attackers to execute backdoor attacks, fine-tuning the adversarial LoRA poses challenges in terms of computational power consumption. Wang and Shu, (2023) propose a backdoor activation attack algorithm, named TA2, which does not require fine-tuning. This algorithm first generates steering vectors by calculating the differences in activations between the clean output and the output produced by a non-aligned LLM. TA2 determines the most effective intervention layer through comparative search and incorporates the steering vectors into the feedforward network. Finally, the steering vectors manipulate the responses of LLMs during the inference.
@@
Table 6: Special instruction for backdoor attacks.
Zhang et al., (2024) introduce an instruction-based backdoor attack method to explore the security of LLMs. They implant backdoors in LLMs solely through designing prompts with embedded backdoor instructions. By utilizing only malicious instructions and corresponding triggers, without the need for any fine-tuning or modification of the LLM parameters, attackers can successfully manipulate the language model. In this study, triggers of various types, including word-level, syntax-level, and semantic-level, are validated, highlighting the potential vulnerabilities of LLMs.
@@ 5 Brief Discussion on Defending Against Backdoor Attacks
Sample Detection In defending against backdoor attacks, defenders prevent the activation of backdoors in compromised models by identifying and filtering out poisoned samples or triggers (Kurita et al.,, 2020; Fan et al.,, 2021; Sun et al.,, 2023). Qi et al., 2021a propose the ONION algorithm, which detects whether the sample has been implanted with the trigger by calculating the impact of different tokens on the sample’s perplexity. The algorithm effectively counters backdoor attacks based on character-level triggers but struggles to defend against sentence-level and abstract grammatical triggers. Shao et al., (2021) observe the impact of removing words on the model’s prediction confidence, thereby identifying potential triggers. They prevent the activation of backdoors by deleting trigger words and reconstructing the original sample. Yang et al., 2021b calculate the difference in confidence between the original samples and the perturbed samples in the target label to detect poisoned samples. The algorithm significantly reduces computational complexity and saves substantial computational resources. Li et al., 2021c propose the BFClass algorithm, which pre-trains a trigger detector to identify potential sets of triggers. Simultaneously, it utilizes the category-based strategy to purge poisoned samples, preserving the model’s security. Li et al., 2021b combine mixup and shuffle strategies to defend against backdoor attacks, where mixup reconstructs the representation vectors and labels of samples to disrupt triggers, and shuffle alters the order of original samples to generate new ones, further enhancing defense capabilities. Jin et al., (2022) hypothesize that essential words should remain independent of triggers. They first utilize weakly supervised learning to train on reliable samples, and subsequently develop a binary classifier that discriminates between poisoned and reliable samples. Zhai et al., (2023) propose a noise-enhanced contrastive learning algorithm to improve model robustness. The algorithm initially generates noisy training data, and then mitigates the impact of backdoors on model predictions through contrastive learning. Wei et al., (2024) design a poisoned sample detector that identifies poisoned samples based on the prediction differences between the model and its variants.
@@ 6 Discussion and Open Challenges
Many backdoor attacks targeting foundational and large language models have been proposed so far, which are described in detail. However, new challenges pertaining to backdoor attacks are arising incessantly. Therefore, there are still some open issues that deserve to be thoroughly discussed and studied. To this end, we provide detailed suggestions for future research directions below.

### v3

@@ 3.1 Backdoor Attack based on Full-parameter Fine-tuning
The efficiency of LLMs has been proven in various NLP tasks, demonstrating their ability to understand and generate text in ways that are both sophisticated and contextually relevant (Xiao et al.,, 2022, 2024). These models have become indispensable tools in machine translation (Zhang et al.,, 2023; Garcia et al.,, 2023), summary generation (Nguyen et al.,, 2021; Nguyen and Luu,, 2022; Zhao et al.,, 2022; Zhao et al., 2023a, ), and recommendation systems (Ma et al.,, 2016; Li et al., 2024a, ). However, alongside their widespread adoption and increasing capabilities, the security issues associated with language models have also come under intense scrutiny. Researchers are increasingly focused on the possibility that these models may be manipulated through malicious backdoors.
@@
You et al., (2023) introduce a backdoor attack algorithm, named LLMBkd, which leverages LLMs to automatically embed a specified textual style as a trigger within samples. Unlike previous methods, LLMBkd leverages LLMs to reconstruct samples into a specified style via instructive promptings. Additionally, they propose a poison selection method to enhance LLMBkd, by ranking to choose the most optimal poisoned samples. Kandpal et al., (2023) explore the security of LLMs based on in-context learning. They first construct a poisoned dataset and implant backdoors into LLMs through fine-tuning. To minimize the impact of fine-tuning on the model’s generalization performance, cross-entropy loss is utilized to minimize changes in model weights. Although this method achieved a high attack success rate, it compromised the model’s performance in translation tasks.
@@
Shi et al., (2023) construct BadGPT, the first backdoor attack against reinforcement learning fine-tuning in LLMs. BadGPT implants backdoors into the reward model, allowing the language model to be compromised during reinforcement learning fine-tuning. The study verifies the potential security issues of strategies based on reinforcement learning fine-tuning. Wang et al., 2023b explore the potential security issues of RLHF, where attackers manipulate ranking scores by altering the rankings of any malicious text, leading to adversarially guided responses from LLMs. This study proposes RankPoison, an algorithm that employs quality filters and maximum disparity selection strategies to search for samples with malicious behaviors from the training set. Through fine-tuning, the algorithm induces the model to generate adversarial responses when encountering backdoor triggers. Li et al., 2023a utilize black-box generative models, such as ChatGPT, as a backdoor attack tool to construct the BGMAttack algorithm. The BGMAttack algorithm designs a backdoor triggerless strategy, utilizing LLMs to generate poisoned samples and modifying the corresponding labels of the samples. Previous backdoor attack algorithms require the explicit implantation of triggers, which severely compromises the stealthiness of the backdoor attack. Zhao et al., 2023b employ manually written prompt as trigger, obviating the need for implanting additional triggers and preserving the integrity of the training samples, enhancing the stealthiness of the backdoor attack. Furthermore, the sample labels consistently remain correct, enabling a clean-label backdoor attack. Tan et al., (2023) propose a more flexible backdoor attack algorithm, named TARGET, which utilizes GPT-4 as a backdoor attack tool to generate malicious templates that act as triggers. The above method requires attackers to possess task-relevant information, which limits its practicality.
Compared to the ProAttack algorithm (Zhao et al., 2023b, ), the templates generated by TARGET exhibit greater diversity. Qi et al., (2023) validate the fragility of the safety alignment of LLMs across three dimensions. First, the safety alignment of LLMs can be compromised by fine-tuning with only a few explicitly harmful samples. Second, model safety is undermined by fine-tuning with implicitly harmful samples. Finally, under the influence of "catastrophic forgetting" (Kirkpatrick et al.,, 2017; Luo et al., 2023b, ), model safety still significantly deteriorates even when fine-tuning on the original dataset. Li et al., 2024c introduce the BadEdit backdoor attack framework, which directly modifies a small number of LLM parameters to efficiently implement backdoor attacks while preserving model performance. Specifically, the backdoor injection problem is redefined as a knowledge editing problem. Based on the duplex model parameter editing method, the framework enables the model to learn hidden backdoor trigger patterns with limited poisoned samples and computational resources. Zou et al., (2024) explore the security of retrieval-augmented generation (RAG) in LLMs. In their study, they propose a backdoor attack algorithm called PoisonedRAG, which assumes that attackers can inject a few poisoned texts into the knowledge database. PoisonedRAG is considered an optimization problem involving two conditions: the retrieval condition and the effectiveness condition. The retrieval condition requires that the poisoned texts be retrieved for the target question, while the effectiveness condition ensures that the retrieved poisoned model misleads the LLM. Yang et al., (2024) investigate the security of LLM-based agents when faced with backdoor attacks. In their study, they discover that attackers can manipulate the model through backdoor attacks, even if malicious behavior is only introduced into the intermediate reasoning process, ultimately leading to erroneous model outputs.
@@
Trends and Challenges Existing work has demonstrated that language models are susceptible to manipulation through backdoors. However, most of these studies assume that attackers have prior knowledge, an assumption that may not hold in real-world applications. Therefore, the following are some trends and challenges in backdoor attacks:
  • Exploring task-agnostic backdoor attack algorithms, which are more challenging and represent a trend that deserves continuous scrutiny.
  • The full-parameter fine-tuning strategy also introduces additional overhead to the deployment of backdoor attacks, which significantly increases the complexity of implementing backdoor attacks.
  • Avoiding the full-parameter fine-tuning of LLMs for the deployment of backdoor attacks, which helps maintain the models’ generalizability, has emerged as a prevalent trend.
@@ 3.2 Backdoor Attack based on Parameter-Efficient Fine-Tuning
Figure 4: Overview of the backdoor attack based on PEFT, where the fine-tuning algorithm employs prompt-tuning. The upper part of the figure illustrates a normal model fine-tuned based on PEFT, while the lower part shows a victim model embedded with backdoors during the fine-tuning process.
@@
Prompt-tuning Xue et al., (2024) introduce TrojLLM, a black-box framework that includes the trigger discovery algorithm and the progressive Trojan poisoning algorithm, capable of autonomously generating triggers with universality and stealthiness. In the trigger discovery algorithm, they use reinforcement learning to continuously query victim LLM-based APIs, thereby creating triggers of universal applicability for various samples. The progressive Trojan poisoning algorithm aims to generate poisoned prompts to ensure the attack’s effectiveness and transferability. Yao et al., (2024) introduce a novel two-stage optimization backdoor attack algorithm that successfully compromises both hard and soft prompt-based LLMs. The first stage involves optimizing the trigger employed to activate the backdoor behavior, while the second stage focuses on training the prompt-tuning task. Huang et al., 2023a propose a composite backdoor attack algorithm with enhanced stealth, named CBA. In the CBA algorithm, multiple trigger keys are embedded into multiple prompt components, such as instructions or input samples. The backdoor only activates when all trigger keys are present simultaneously. This algorithm balances anomaly strength in the prompt and minimizes semantic changes, which is more effective than simple combinations of triggers (Yang et al., 2021c, ).
@@
Trends and Challenges The effectiveness of backdoor attacks, particularly those that target PEFT methods, has been clearly demonstrated. Below are some trends and challenges in backdoor attacks based on parameter-efficient fine-tuning algorithms:
  • Existing work primarily focuses on classification tasks; however, a new trend is exploring backdoor attacks targeting generative tasks, such as question-answering or knowledge reasoning.
  • Unlike classification tasks, backdoor attack algorithms targeting generation tasks often require malicious modification of sample labels. Although these modifications can achieve effective attack results, they may compromise the stealthiness of backdoor attack. Therefore, exploring clean-label backdoor attacks in generation tasks presents a significant challenge.
@@ 3.3 Backdoor Attack without Fine-tuning
Figure 5: Overview of the backdoor attack without fine-tuning, illustrating attacks on instructions and in-context learning. Attackers manipulate model responses through malicious instructions and poisoned demonstration examples.
Wang and Shu, (2023) propose a backdoor activation attack algorithm, named TA2, which does not require fine-tuning. This algorithm first generates steering vectors by calculating the differences in activations between the clean output and the output produced by a non-aligned LLM. TA2 determines the most effective intervention layer through comparative search and incorporates the steering vectors into the feedforward network. Finally, the steering vectors manipulate the responses of LLMs during the inference.
LoRA [^2] It is noteworthy that although LoRA is employed, the subsequent work, as an exception, does not specifically fine-tune LLMs for backdoor attacks. In share-and-play settings, Liu et al., 2024a assume that the LoRA (Hu et al.,, 2021) algorithm could be a potential attacker capable of injecting backdoors into LLMs. They combine an adversarial LoRA with a benign LoRA to investigate attack methods that do not require backdoor fine-tuning. Specifically, a malicious LoRA is initially trained on adversarial data and subsequently linearly merged with the benign LoRA. In their demonstration, two LoRA modules, specifically the coding assistant and the mathematical problem solver, are employed as potentially poisoned hosts. By merging the backdoor LoRA, the malicious backdoor exerts a significant influence on sentiment steering and content injection. Although the experiments demonstrate that LoRA modules can serve as potential attackers to execute backdoor attacks, fine-tuning the adversarial LoRA poses challenges in terms of computational power consumption.
@@
Table 4: Special malicious instruction for backdoor attacks.
Zhang et al., (2024) introduce an instruction-based backdoor attack method to explore the security of LLMs. They implant backdoors in LLMs solely through designing prompts with embedded backdoor instructions. By utilizing only malicious instructions and corresponding triggers, without the need for any fine-tuning or modification of the LLM parameters, attackers can successfully manipulate the language model. In this study, triggers of various types, including word-level, syntax-level, and semantic-level, are validated, highlighting the potential vulnerabilities of LLMs. The method described above requires attackers to obtain deployment permissions for the model, a challenging requirement in most applications.
Trends and Challenges It has been proven that attackers can manipulate model responses merely through malicious instructions or poisoned demonstration examples. Some new challenges and trends need attention:
  • Although this paper only considers backdoor attacks targeting large language models in NLP, exploring backdoor attacks without fine-tuning in large vision-language models (Liang et al.,, 2024) or multimodal decision systems (Jiao et al.,, 2024) is an emerging trend.
  • Backdoor attacks based on malicious instructions (Wang et al., 2023a, ) and poisoned demonstration examples (Zhao et al., 2024b, ) have proven to be effective. However, their explicit triggers are easily recognized by defense algorithms. Consequently, exploring more covert triggers in backdoor attacks without fine-tuning represents a challenge that warrants sustained attention.
@@ 5 Brief Discussion on Defending Against Backdoor Attacks
Sample Detection In defending against backdoor attacks, defenders prevent the activation of backdoors in compromised models by identifying and filtering out poisoned samples or triggers (Kurita et al.,, 2020; Fan et al.,, 2021; Sun et al.,, 2023; Zeng et al.,, 2024; Zhao et al., 2024d, ; Liu et al., 2024b, ). Qi et al., 2021a propose the ONION algorithm, which detects whether the sample has been implanted with the trigger by calculating the impact of different tokens on the sample’s perplexity. The algorithm effectively counters backdoor attacks based on character-level triggers but struggles to defend against sentence-level and abstract grammatical triggers. Shao et al., (2021) observe the impact of removing words on the model’s prediction confidence, thereby identifying potential triggers. They prevent the activation of backdoors by deleting trigger words and reconstructing the original sample. Yang et al., 2021b calculate the difference in confidence between the original samples and the perturbed samples in the target label to detect poisoned samples. The algorithm significantly reduces computational complexity and saves substantial computational resources. Li et al., 2021c propose the BFClass algorithm, which pre-trains a trigger detector to identify potential sets of triggers. Simultaneously, it utilizes the category-based strategy to purge poisoned samples, preserving the model’s security. Li et al., 2021b combine mixup and shuffle strategies to defend against backdoor attacks, where mixup reconstructs the representation vectors and labels of samples to disrupt triggers, and shuffle alters the order of original samples to generate new ones, further enhancing defense capabilities. Jin et al., (2022) hypothesize that essential words should remain independent of triggers. They first utilize weakly supervised learning to train on reliable samples, and subsequently develop a binary classifier that discriminates between poisoned and reliable samples. Zhai et al., (2023) propose a noise-enhanced contrastive learning algorithm to improve model robustness. The algorithm initially generates noisy training data, and then mitigates the impact of backdoors on model predictions through contrastive learning. Wei et al., (2024) design a poisoned sample detector that identifies poisoned samples based on the prediction differences between the model and its variants. To mitigate backdoor attacks, the CLEANGEN model Li et al., 2024d replaces suspicious tokens with those generated by the clean reference model. Li et al., 2024b propose a Chain-of-Scrutiny approach, which utilizes demonstrations to guide large language models in generating detailed reasoning steps, ensuring that the model responses align with the final output.
@@
Trends and Challenges Defending against backdoor attacks is crucial for establishing a secure and reliable NLP community, and several new issues merit attention:
  • Traditional defense algorithms predominantly focus on identifying poisoned samples or modifications to the weights of victim models. However, scrutinizing instructions or demonstration examples for potential security vulnerabilities warrants further attention.
  • Similar to backdoor attacks that operate without fine-tuning, the exploration of defense algorithms that also eschew model fine-tuning is worthwhile, significantly augmenting the usability of these mechanisms.
@@ 6 Discussion and Open Challenges
Many backdoor attacks targeting foundational and large language models have been proposed so far, which are described in detail. However, new challenges pertaining to backdoor attacks are arising incessantly. Therefore, there are still some open issues that deserve to be thoroughly discussed and studied, as shown in Figure 6. To this end, we provide detailed suggestions for future research directions below.
{forest}
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Open
Challenges [Backdoor
Attack [Trigger Design
Section 6.1[ How to design more covert and universal backdoor
attack triggers?]] [Clean-labe
Section 6.2[How to deploy clean-label backdoor attacks on
generation tasks?]] [Without Fine-tuning
Section 6.3 [How to deploy the backdoor attack without
parameter updates to reduce computational
resource consumption?]] ] [Defend Against
Backdoor Attack [Defenses
Section 6.4 [How to design a high-performance and universal
defense algorithm?]] [Backdoor Evaluation
Section 6.5 [How to proactively check if a model has
been compromised?]] ] [Others
Section 6.6 [Interpretation Analysis; Evaluation Metrics; Uniform Benchmark]] ]
Figure 6: Open Challenges in Backdoor Attacks on Large Language Models.